the perimeter of a rectangle is 64ft. describe the possible lengths of a side if the area of the rectangle is not to exceed 192 square feet.
Answer by sravanik2005(14)

perimeter,2(l+w)=64 => l+w=32
and lw<=192
l(32-l)<=192
l^2-32l+192<=0
(l-24)(l-8)<=0
8<=l<=24
and 8<=w<=24
hence possible lengths are (8,24)(9,23)(10,22)(11,21)(12,20)(13,19)(14,18)(15,17)(16,16)